# Mechanistic Unlearning: Robust Knowledge Unlearning and Editing via Mechanistic Localization

### v1

@@ 1 Introduction
However, the effectiveness of these “output-tracing” (OT) techniques for editing has been questioned by Hase et al. (2023). Our research confirms these doubts, finding that localized editing and unlearning of facts based on several existing OT methods often perform equal to or worse than simply updating the entire model. This is particularly evident when evaluating the robustness of edits against prompt variations and relearning, and when probing for remaining latent knowledge.
@@
  • We perform a rigorous evaluation of several standard unlearning approaches on factual recall tasks and show they fail to generalize to prompting/output distribution shifts and adversarial relearning.
  • We identify mechanisms for factual lookup and attribute extraction on Gemma-7B and Gemma-2-9B. We demonstrate that gradient-based unlearning and editing localized on the factual lookup mechanism is more robust and generalizes better than OT localizations and baselines across multiple datasets and models.
  • We analyze intermediate representations using probing, and provide further evidence that editing with FLU localization reduces the available latent information more than other localizations and baselines. We also analyze the weights that are modified for each localization, and find that OT techniques and baselines modify the attribute extraction mechanisms more than the fact lookup mechanism.
@@ 1.1 Related Work
A growing body of work aims to unlearn a subset of the training data in LLMs. Eldan & Russinovich (2023) propose a method for unlearning entire books like the Harry Potter series. Chen & Yang (2023) consider modifying transformer architecture by inserting “unlearning” layers.
@@
In the context of LLMs and safety, techniques such as Helpful-Harmless RLHF (Bai et al., 2022) and Representation Misdirection for Unlearning (Li et al., 2024) aim to suppress dangerous knowledge or harmful tendencies in LLMs. A related line of work on safety proposes methods making it difficult to modify open models for use on harmful domains (Deng et al., 2024; Henderson et al., 2023).
@@
have been shown for both localized and nonlocalized methods. Patil et al. (2023) extract correct answers to edited facts from the intermediate residual stream and through prompt rephrasing. Yong et al. (2024) show that low-resource languages jailbreak models output unsafe content, and Lo et al. (2024); Lermen et al. (2023) demonstrate that relearning with a small amount of compute/data causes models to regain undesirable knowledge/tendencies. Even without explicit finetuning, Xhonneux et al. (2024) show that in-context learning alone suffices to reintroduce undesirable knowledge despite the model being designed to refuse to output such knowledge. Lee et al. (2024) shows that even after alignment techniques are applied to make models nontoxic, toxicity representations are still present, just not triggered - they argue that this is a reason that models lack robustness and can still be jailbroken to trigger this unwanted behavior.
@@ 2 Methods
Our experiments are designed to test the effectiveness of localization for unlearning/editing of facts. In this section we describe the tasks used and the localization and unlearning methods evaluated.
2.1 Unlearning/Editing Tasks
We focus on unlearning and editing subsets of two datasets: (1) Sports Facts dataset from Nanda et al. (2023), which contains subject-sport relations across three sports categories for 1567 athletes, and (2) the CounterFact dataset from Meng et al. (2023).
Sports Facts
In the Sports Facts dataset, we attempt to unlearn and edit two groups of factual associations. First, we unlearn all athlete-sport associations for a given sport. In this case, we establish a forget set consisting of all the basketball athletes. Second, we edit factual associations for a set of 16 athletes belonging to all three sports categories. We test editing these sets of associations by replacing their sports with golf, and the retain set is the rest of the non-forget athletes. We use the Gemma-7B LLM (Team et al., 2024) rather than the Pythia-2.8B (Mallen & Belrose, 2023) model tested in Nanda et al. (2023), for its stronger general capabilities which we can measure for side effects, and for its ability to provide sports knowledge in different input/output formats.
CounterFact
In the CounterFact dataset, following Geva et al. (2023), we first filter the dataset for facts which the model assigns higher than 50% probability to the right answer, leaving 2170 facts. Then, we edit a set of 16 facts, replacing the correct answers with an alternative false target, with the retain set being the rest of the non-forget facts. We use the Gemma-2-9B LLM, as an alternative to the Gemma-7B LLM, to test a range of model architectures and sizes. We don’t test smaller models as they cannot generally output knowledge in different formats, as measured by our robustness evaluations.
@@ 2.2 Localization Methods and Baselines
Given a model $M:X\mapsto L$ mapping sequence of tokens $X$ to logits $L\in\mathbb{R}^{|V|}$ over vocabulary $V$, we consider $M$ to be a directed acyclic graph $(C,E)$ with $C$ being a set of model components and $E$ being edges between components. Adopting notation from Elhage et al. (2021), we consider the query, key, value, and output weights $W_{Q}^{h},W_{K}^{h},W_{V}^{h},W_{O}^{h}$ of each head along with the input and output projection weights $W_{I}^{m},W_{O}^{m}$ of each MLP as components.
@@
First, we test Causal Tracing, a method for finding components with high direct causal importance for factual associations (Meng et al., 2023). Previous work has highlighted the shortcomings of Causal Tracing as a localization method (Hase et al., 2023), so we also use Attribution Patching (Nanda, 2023), which uses a linear approximation to activation patching to automatically localize over components with high direct and indirect importance.
@@
We hypothesize that these output-based techniques will prioritize the shared extraction components and other mechanisms for reformatting predictions over the more diffuse FLU components, and thus appear more precise yet leave the underlying latent information present in the model. This might decrease robustness under alternative extraction methods, thus motivating non-OT-based localization, described next.
Fact Lookup (FLU) localization: Manual Mechanistic Interpretability
@@
Next, we use manually derived localizations for MLP layers. For Sports Facts, our localization is inspired by Nanda et al. (2023), who discovered components in Pythia 2.8B responsible for token concatenation, fact lookup, and attribute extraction. They, along with Geva et al. (2023), find that the fact-lookup stage enriches the latent stream with information about the subject (athlete) at the subject’s token position, and the attribute extraction stage extracts the latent sport information and formats it in the final token position. We replicate a key result of their work in Gemma-7B, localizing the fact lookup stage to be the MLP components between layers 2 and 7.
For CounterFact, we replicate findings from Geva et al. (2023) and do further causal analyses to identify particular MLPs that perform fact lookup of our forget set in Gemma-2-9B. Our analysis highlights layers 3-5, 7-10, and 14-17 as the critical MLPs. The manual analysis for both datasets is outlined in Section A.2.1. We also locate causally important attribute extraction mechanisms, both attention heads and MLPs, in later layers of the model.
@@
We refer to these FLU localizations as manual interpretability. Importantly, our analysis differs from OT techniques because we consider the causal effects of ablations upon intermediate representations used by the factual recall mechanism, not just the effects on the output. We hypothesize that the optimal location for robust unlearning/editing is in the fact lookup stage rather than in the attribute extraction stage, because adversaries can develop alternative methods for extracting knowledge from the latent stream through alternative prompts or white-box methods so we want to prevent the knowledge from ever being added to the latent stream. Thus, we exclusively modify the fact lookup MLPs.
Baselines: Random, All-MLPs, and Nonlocalized
@@
We additionally consider three baselines: one corresponding to $C_{\tau}=C$ (i.e., no localization, optimizing all the components of the model), another that randomly chooses components, and another that trains all MLP components. We test the last All-MLPs localization to determine if our mechanistically localized MLPs are uniquely important - we want to know if the same unlearning performance can be achieved with just the heuristic that training only MLPs improves robustness, or if mechanistic understanding of the role of the component is crucial.
@@
In Section A.2.3, we analyze the proportions of each mechanism (the extraction heads, extraction MLPs, and fact lookup MLPs, by parameter count) that are present in each localization.
2.3 Parameter Update Methods
@@
Once we have a localization $C_{\tau}$, we run one of the unlearning or editing methods, restricting weight updates to only components in $C_{\tau}$. We modify weights using gradient ascent/descent on a combination of loss functions. As an alternative method, we also test masking a fraction of the weights as an alternative method showing precision of edits, with results reported in Section 3.2.
@@
Following work by Lee et al. (2023) and Panigrahi et al. (2023), we fine-tune the parameters within the localized components. We use a loss function

### v2

@@ 1 Introduction
We coin a new term to categorize localizations which measure causal effects of components on only the output: Output-Tracing localizations. The effectiveness of output-tracing (OT) techniques like Causal Tracing for editing has been questioned by Hase et al. (2023). Our research confirms these doubts, finding that localized editing and unlearning of facts based on several existing OT methods often perform equal to or worse than simply updating the entire model. This is particularly evident when evaluating the robustness of edits against prompt variations and relearning, and when probing for remaining latent knowledge.
@@
  • We perform a rigorous evaluation of several standard editing approaches on factual recall tasks, and we identify mechanisms for factual lookup and attribute extraction on Gemma-7B, Gemma-2-9B, and Llama-3-8b. We demonstrate that gradient-based editing localized on the factual lookup mechanism is more robust than OT localizations and baselines across multiple datasets, models, and evaluations.
  • We demonstrate that it is more difficult to elicit the forgotten ground truth answers using alternative prompting with FLU localizations. We also demonstrate slower or no relearning of the ground truth answers, retraining edited models on half of the edited set and evaluating them on the other half of the edited set.
  • We analyze intermediate representations using probing, and provide further evidence that editing with FLU localization modifies the internal latent information to reflect the desired edited answer more than other localizations and baselines. We also analyze the weights that are modified for each localization, and find that OT techniques and baselines modify the attribute extraction mechanisms more than the fact lookup mechanism.
@@ 1.1 Related Work
In the context of LLMs and safety, techniques such as Helpful-Harmless RLHF (Bai et al., 2022) and Representation Misdirection for Unlearning (Li et al., 2024b) aim to suppress dangerous knowledge or harmful tendencies in LLMs. Li et al. (2024a); Zou et al. (2023; 2024) approach unlearning and dangerous knowledge suppression from a top-down feature view, reading or suppressing linear features related to memorized, harmful, and undesired concepts. A related line of work on safety proposes methods making it difficult to modify open models for use on harmful domains (Tamirisa et al., 2024; Deng et al., 2024; Henderson et al., 2023), including through adversarial relearning.
@@
have been shown for both localized and nonlocalized methods. Patil et al. (2023) extract correct answers to edited facts from the intermediate residual stream and through prompt rephrasing. Yong et al. (2024) show that low-resource languages jailbreak models output unsafe content, and Lo et al. (2024); Lermen et al. (2023); Deeb & Roger (2024) demonstrate that relearning with a small amount of compute/data causes models to regain undesirable knowledge/tendencies. Even without explicit finetuning, Xhonneux et al. (2024) show that in-context learning alone suffices to reintroduce undesirable knowledge despite the model being designed to refuse to output such knowledge. Lee et al. (2024) shows that even after alignment techniques are applied to make models nontoxic, toxicity representations are still present, just not triggered - they argue that this is a reason that models lack robustness and can still be jailbroken to trigger this unwanted behavior.
Hong et al. (2024) evaluates unlearning by measuring residual knowledge left in internal activations, and demonstrate that current approaches fail to remove this residual knowledge and thus can be exploited. They attempt unlearning by targeting the MLPs these residual knowledge traces reside in, but fail to find a non-oracle unlearning approach that successfully removes residual information.
@@ 2 Methods
Our experiments are designed to test the effectiveness of localization for editing of facts. In this section we describe the tasks used and the localization and editing methods evaluated.
2.1 Editing Tasks
We focus on editing subsets of two datasets: (1) Sports Facts dataset from Nanda et al. (2023), which contains subject-sport relations across three sports categories for 1567 athletes, and (2) the CounterFact dataset from Meng et al. (2023).
Sports Facts: Sports-Athlete-Editing, Full-Sports-Editing, and Sports-Unlearning tasks
In the Sports Facts dataset, we edit two general groups of factual associations. For the first editing task, we edit factual associations for a constant set of randomly selected athletes belonging to any of the three sport categories. We test editing these sets of associations by replacing their correct sports with one of the other two incorrect sports (with equal probability). To increase the comprehensiveness of our evaluation, we run experiments with different forget set sizes: 16 athletes and 64 athletes. We refer to this task as Sports-Athlete-Editing. These chosen forget sets are constant between all localizations. For the second editing task, we unlearn all athlete-sport associations for athletes who play one sport. In this case, we establish a forget set consisting of all the athletes who play one sport (basketball, baseball, or football), and we edit the association by replacing the athlete’s correct sport with golf. For comprehensiveness, we vary the sport that the forget set is constructed from. We refer to this task as Full-Sports-Editing. Finally, we also design an unlearning task, Sports-Unlearning, where the goal is to unlearn factual associations for all athletes in one of the sports.
CounterFact-Editing and Sequential-CounterFact-Editing task
In the CounterFact dataset, following Geva et al. (2023), we first filter the dataset for facts which our models assign higher than 50% probability to the right answer, which varies per model. The goal of our CounterFact-Editing task is to edit a constant set of facts, replacing the correct answers with an alternative false target, with the retain set being the rest of the non-forget facts. We vary forget set sizes to be of 16 and 64 facts. In Sequential-CounterFact-Editing task, we edit a total of 64 facts by sequentially editing four randomly selected subsets of 16 facts. We test sequential editing here because facts from CounterFact can reside in different parts of the model, so we wish to test if we can exploit different localizations for different facts. These chosen forget sets are constant between all localizations.
Models
We implement editing on the Gemma-7B LLM, the Gemma-2-9B LLM, and the Llama-3-8b LLM. We don’t use the Pythia-2.8B (Biderman et al., 2023) and GPT-2 models tested in the previous fact interpretability literature because our larger models have stronger general capabilities which we can measure for side effects, and also because our larger models can provide factual knowledge in more input/output formats for more robustness evaluations. All graphs presented in the main text and appendix, unless otherwise specified, are averaged over all three model types.
@@ 2.2 Localization Methods and Baselines
Given a model $M:X\mapsto L$ mapping sequence of tokens $X$ to logits $L\in\mathbb{R}^{|V|}$ over vocabulary $V$, we consider $M$ to be a directed acyclic graph $(C,E)$ with $C$ being a set of model components and $E$ being edges between components. Adopting notation from Elhage et al. (2021), we consider the query, key, value, and output weights of each head along with the input, gate, and output projection weights of each MLP as components.
@@
First, we test Causal Tracing, a method for finding components with high direct causal importance for factual associations (Meng et al., 2023). We also use Attribution Patching (Nanda, 2023) as a fast and acceptably accurate approximation of causal tracing to automatically localize over components with high direct and indirect importance. We additionally consider the versions of these localizations with only MLPs (Causal Tracing MLPs and Attribution Patching MLPs).
@@
We hypothesize that these output-based techniques will prioritize the shared extraction components and other mechanisms for reformatting predictions over the more diffuse FLU components, and thus appear more precise yet leave the underlying latent information present in the model. This might decrease robustness under alternative extraction methods, thus motivating non-OT-based localization, described next. We discuss the precise components/mechanisms highlighted by OT localizations in Section A.3.
Fact Lookup (FLU) localization:
@@
Next, we use manually derived localizations for MLP layers. For Sports Facts, our localization is inspired by Nanda et al. (2023), who identified components in Pythia 2.8B responsible for token concatenation, fact lookup, and attribute extraction. Their work, along with Geva et al. (2023), demonstrates that the fact-lookup stage enriches the latent representation of the subject (athlete) with information about their corresponding sport, while the attribute extraction stage extracts the latent sport information and formats it in the final token position. We replicate a key result of their work in our three models by training logistic regression models (probes) to predict the correct sport using the internal activations of the model taken from different layers. We consider the FLU localization to be the layers at which the probe accuracies rapidly increase as these correspond to layers where representations of the athletes are being enriched to encode the correct sport.
For CounterFact, we cannot use the probe technique since unlike Sports Facts, the correct answers for the dataset do not fall into a few categories that we can train a probe for. Instead, we first use path patching from Goldowsky-Dill et al. (2023) to measure the direct importance of attention heads and MLPs on the final logit difference between the correct answer and a sample incorrect answer. Path patching is a technique that finds the effect of corrupting a single path from a sender component to a receiver component on the final logits of the model. Components that change the output significantly without being mediated by other components do so by directly affecting the logits, and thus we consider them to be responsible for extracting the facts encoded in the representation into the answer logits. We thus refer to these components as the “attribute extraction mechanism” Geva et al. (2023). We use path patching again, this time patching paths between MLPs and the attribute extraction mechanism to find the components with the highest contribution to the logit difference as mediated through the extraction mechanisms. Such components enrich the token representations with the appropriate facts to then be extracted, and thus we use them as our FLU localization. More details about the manual analysis for both datasets is outlined in Section A.2.1.
@@
Importantly, FLU differs from OT techniques by focusing on the causal effects of ablations upon intermediate representations used by the factual recall mechanism, not just the effects on the output. We hypothesize that the optimal location for robust editing is in the fact lookup stage rather than in the attribute extraction stage,as adversaries could potentially devise alternative extraction methods if knowledge remains in the latent stream. Therefore, our model edits are focused exclusively on the fact lookup MLPs.
Baselines: Random-MLPs, Random, All-MLPs, and Nonlocalized
@@
We additionally consider four baselines: one corresponding to $C_{\tau}=C$ (i.e., no localization, optimizing all the components of the model), another that randomly chooses components, another that trains all MLP components, and another trains a random selection of MLP components. We test the last MLP baselines to determine if our mechanistically localized MLPs are uniquely important - we want to know if the same unlearning performance can be achieved with just the heuristic that training only MLPs improves robustness, or if mechanistic understanding of the role of the component is crucial.
@@
In Section A.5, we analyze the proportions of each mechanism (the extraction heads, extraction MLPs, and fact lookup MLPs, by parameter count) that are present in each localization.
The main text focuses on comparing FLU, Causal Tracing, and Nonlocalized, while the appendix has the same figures with all above localizations included, with the same conclusions in every case.
2.3 Parameter Update Method
@@
Once we have a localization $C_{\tau}$, we run one of the unlearning or editing methods, restricting weight updates to only components in $C_{\tau}$. We update weights using gradient descent on a combination of loss functions.
@@
Following work by Lee et al. (2023) and Panigrahi et al. (2023), we fine-tune the parameters within the localized components. For editing, we use a loss function $L=\lambda_{1}L_{\text{injection}}+\lambda_{2}L_{\text{retain}}+\lambda_{3}L_{\text{SFT}}$, where $L_{\text{injection}}$ is a cross-entropy loss on the forget facts maximizing the probability of the alternative new false target. $L_{\text{retain}}$ is a cross-entropy loss on a train split of the remaining facts, and $L_{\text{SFT}}$ is a cross-entropy loss on the Pile dataset (Gao et al., 2020). We sweep over learning rates and injection loss $\lambda$s for three representative localizations in Section A.6 .
3 Editing Evaluation
In this section, we show the results of model editing with localization: we test localization techniques from Section 2.2, and edit these localized components using fine-tuning (Section 2.3). Here we focus on four main editing tasks: sports-athlete-editing, full-sports-editing, counterfact-editing, and sequential-counterfact-editing. We present augmenting results for sports-unlearning in Section A.1.
All the editing tasks are assessed based on prompt-completion based and adversarial relearning evaluations.
3.1 Prompting-based Evaluation
Our prompting-based evaluation assesses an editing method’s ability to forget or edit specific information while retaining unrelated knowledge. This is measured by evaluating how the model post-editing completes the prompts coming from the forget set: we report how accurately it recalls the undesired forgotten answer (forget accuracy), and how accurately it recalls the new desired edited answer (edit accuracy). In addition, we also measure the accuracy on facts not in the forget set (maintain accuracy). In cases where a positive result is lower accuracy, we use the term error to denote 1 - accuracy (e.g. forget error = 1 - forget accuracy). Thus, well edited models should decrease forget accuracy, increase forget error, and increase edit accuracy. Results of these standard evaluations are reported in Section A.7.1.
Inspired by Patil et al. (2023) and Lynch et al. (2024), to ensure the editing process has not overfit to the specific format of the original prompts, we incorporate a robustness check using a multiple-choice question format (MCQ accuracy). This helps determine to what extent the model edited the information, and whether it can still access and utilize that knowledge when prompted differently. In this MCQ evaluation, the prompt also includes some in-context examples of answering multiple choice questions correctly. On the forget set, we refer to the accuracy of the model answering with the ground truth as the MCQ Forget Accuracy (stronger methods should decrease MCQ Forget Accuracy), and the accuracy of the model answering with the new edited answer’s choice as the MCQ Edit Accuracy (stronger methods should increase MCQ Edit Accuracy).
Finally, we also evaluate the models’ accuracy post-editing on MMLU (Hendrycks et al., 2021) as a proxy for general language understanding to measure any unintended side effects.
3.1.1 Sports Tasks
Sports Prompting
For the sports dataset, following Nanda et al. (2023), we first evaluate the accuracy of our models to complete the prompt, “Fact: [athlete] plays the sport of”, with a one-shot example of Tiger Woods playing golf given first. Note that this is the same prompt used for the editing loss in the first place. For the MCQ evaluation, we use choices of all four sports (football, baseball, basketball, and golf).
We average accuracies over all models, for Sports-Athlete-Editing we average over editing both 16 and 64 facts, and for Full-Sports-Editing we average over editing Basketball, Baseball, and Football.
Sports Results
(a) Sports-Athlete-Editing
(b) Full-Sports-Editing
Figure 2: Spider plots illustrating the advantages of FLU for editing Sports across adversarial prompting and relearning evaluations, averaged over all three model types. (Left) The Sports-Athlete-Editing plot shows that FLU localization leads to editing that is the most robust against MCQ prompting and relearning. (Right) The plot shows that most localizations perform approximately equivalently in the Full-Sports-Editing task, with FLU localization slightly better for MCQ.
Our analysis reveals that editing employing FLU localization exhibits superior performance in forgetting the original information and adopting the edited information, across different prompt formats. As explained in Section 3.1, better editing should result in higher MCQ Forget Error, and higher MCQ Edit Accuracy. Figure 3 shows that FLU localized models are the best on both fronts. The difference is especially significant in Figure 3a, where only FLU model edits generalize meaningfully to MCQ, exceeding other localization methods by more than 40% in MCQ Edit Accuracy.
A comprehensive comparison of all localization methods with multiple-choice prompting is available in Section A.7.2, further supporting our findings.
(a) Sports-Athlete-Editing
(b) Full-Sports-Editing
Figure 3: Bar charts showing results of MCQ evaluations, reporting both the forget error and edit accuracy when prompted with MCQ, averaged over three model types. For both (a) Sports-Athlete-Editing and (b) Full-Sports-Editing, FLU localization answers with the original answer the least (MCQ Forget Error) and answers with the edited answer most accurately (MCQ Edit Accuracy).
3.1.2 CounterFact Tasks
CounterFact Prompting
For CounterFact, we create an MCQ evaluation with four choices for every question, randomly ordering the true answer, the injected false answer, and two other question-specific LLM-generated incorrect answers. We also consider the original robustness and side effect evaluations from the Meng et al. (2023) dataset: the Paraphrase and Neighborhood facts accuracies. Answers of edited facts are meant to generalize to the Paraphrase evaluation, which phrases the fact in a different but equivalent way, so we report the Paraphrase Edit Accuracy (stronger methods should increase Paraphrase edit accuracy). Editing should not generalize to the Neighborhood evaluation, which presents similar but unrelated facts to the forget set, so we report the Neighborhood Edit Error which is lower if models incorrectly report the edited answer in these unrelated facts.
We again also use an MCQ evaluation, where the choices consist of the true answer, the injected false answer, and two other question-specific LLM-generated incorrect answers. We note that Paraphrase and Neighborhood evaluations ask for the answer in the same original format, so they are more in-distribution than MCQ. We average accuracies over all models, and for CounterFact-Editing we average over editing both 16 and 64 facts.
CounterFact Results
Figure 5 illustrates the robustness of FLU editing in the MCQ and Paraphrase evaluations. Edited models using FLU localization answer less frequently with the original, incorrect information (MCQ Forget Error) and more frequently provide the new, edited answer. Furthermore, the Neighborhood Edit Error highlights that other localization methods exhibit slightly more pronounced side effects, inadvertently editing unintended, semantically similar facts.
Interestingly, sequential editing displays marginally greater robustness than nonsequential editing in MCQ when comparing between CounterFact-Editing and Sequential-CounterFact-Editing bars. This observation supports an approach that editing large sets of facts can be made more effective by partitioning the set and applying edits sequentially. We present results comparing all localizations across each prompt robustness evaluation in Section A.7.2, with the largely consistent conclusions.
(a) CounterFact-Editing
(b) Sequential-CounterFact-Editing
Figure 4: Spider plots illustrating the advantages of FLU for editing CounterFact across prompting evaluations, averaged over all three model types. (Left) The CounterFact-Editing plot shows that FLU localization leads to editing that is the most robust against MCQ prompting and Paraphrasing. (Right) The Sequential-CounterFact-Editing plot shows that FLU localization is the most robust against MCQ prompting.
(a) CounterFact-Editing
(b) Sequential-CounterFact-Editing
Figure 5: Bar charts showing results of MCQ, Paraphrase, and Neighborhood prompt evaluations, averaged over all three model types. For both (a) CounterFact-Editing and (b) Sequential-CounterFact-Editing, FLU localization has the most robust edit accuracy measured by MCQ and Paraphrase. FLU localization editing also does not incorrectly generalize to Neighborhood prompts. Sequential editing is slightly more robust than nonsequential editing in MCQ when comparing between CounterFact-Editing and Sequential-CounterFact-Editing bars.
3.2 Adversarial Relearning Evaluation
We measure the ability of our models to withstand adversarial relearning, both to address the scenario in which adversaries may have fine-tuning access and as a measure for the quality of editing. We replicate the methodology of Deeb & Roger (2024), splitting our forget sets in two independent halves, retraining with half of the ground truth labels, and evaluating on the other half. This methodology aims to discern whether the editing technique successfully removed the underlying factual association or merely obfuscated its direct retrieval while leaving it potentially susceptible to recovery when doing partial retraining.
We retrain with a rank-512 LoRA across all linear modules, with details available in Section A.7.3. In this section, we focus on the Sports-Athlete-Editing task, as in the other tasks it was either too easy to relearn (Full-Sports-Editing) or too hard to relearn any performance (all CounterFact tasks) across all localizations. Relearning isn’t a valid evaluation for Full-Sports-Editing because the facts are not independent, and models should reasonably generalize from relearning on half the basketball athletes to answering with basketball on the other half of the basketball athletes. We show results from Counterfact and from all localizations in Section A.7.3.
Sports Results
Our adversarial relearning experiments, as depicted in Figure 6, reveal that retraining on a subset of the original “forgotten” data can recover a significant portion, as much as 63%, of the supposedly forgotten information when using OT methods like Causal Tracing and Causal Tracing MLPs. This suggests that these methods may simply mask direct retrieval of this information, leaving the model susceptible to this information recovery through retraining. In contrast, FLU localization exhibits greater resilience to such adversarial relearning, with only about 20% of the forgotten information recovered. This indicates that FLU localization may be more effective in targeting and removing the underlying knowledge, making it harder to recover through retraining.
Figure 6: Relearning recovers the least accuracy on the forget set using FLU localizations. Relearning recovers significant accuracy on the original forget set in OT localizations (Causal Tracing and Causal Tracing MLPs).
3.3 Latent Knowledge Analysis
In this section, we provide more evidence of our hypothesis that FLU unlearning targets the source of intermediate latent knowledge. We analyze the Sports-Athlete-Editing task again here because the ground truth and the edited answers vary between one of only three possibilities.
We train logistic regression models (probes) (Alain & Bengio, 2018) on prompt activations following every model layer to predict the correct ground truth sport on the maintained set of athletes. This is possible because there are only three possible sports, so we can train binary classification probes for each sport and take the maximum classification over the sports. This is meant to discover internal representations of the true sport the model believes the answer to be: then, we apply these probes on the activations of the forget set of athletes.
We present graphs averaged over models and over 16 and 64 facts. Because layer is on the x-axis and models have different numbers of layers, averaging performance at a layer over multiple models isn’t perfectly valid, so we also present probing graphs for each individual model and all localizations in Section A.7.4.
Sports Results
Figure 7: Linear probes applied to the forget set, classifying model activations after various layers. (Left) The line graph shows that some localizations still represent almost completely correct forget set knowledge in early layers, especially OT, while FLU localizations represent this original knowledge the least. (Right) The line graph shows that FLU localizations represent the edited rather than original answer earlier and more consistently throughout layers than any other localization.
In Figure 7, the probes on FLU consistently predict the forget answer less and the edit answer more than in any other localization, especially in early layers. Furthermore, the FLU probe classifications for the most part monotonically converge from their nonzero starting accuracy to 0 (for forget accuracy) and 1 (for edit accuracy).
Every other localization has much higher peak probe classification forget accuracy in the early layers, especially the OT localizations which have peak classification forget accuracy of almost 100%. This strongly suggests that these models still significantly represent the ground truth answer rather than the edit answer in early layers.
3.4 The Role of Parameter Count
In this section, we perform weight-masking to quantify the size of edits with different localizations, as well as to investigate which factual mechanisms are targeted when editing with different localizations. We employ a weight masking technique involving training a binary differentiable mask over individual weights of the model within the localized components, inspired by weight pruning/masking work (Bayazit et al., 2023; Panigrahi et al., 2023). In this case, no weight updates are being performed. Rather, the mask turns a subset of the weights to zero.
Controlling for Parameter Count
Although we already standardize the number of trainable parameters in most localizations, we additionally investigate if FLU editing is better than other localization techniques when controlling for the exact number of parameters that are masked.
We perform weight masking on the Sports-Unlearning, Sports-Athlete-Editing, and CounterFact-Editing tasks. Detailed results are reported in Section A.4. We find that when controlling for the size of the localization FLU is consistently more robust when subject to our suite of evaluations.
Other Localizations Affect the Extraction Mechanism
After training weight masks, we analyze the proportion of each mechanism (fact lookup, attribution extraction) that is masked by each localization’s weight mask in Section A.5. We demonstrate that OT methods and nonlocalized editing all modify a higher proportion of the extraction head/MLP parameters than the fact lookup mechanism parameters, supporting our claim that OT methods target extraction mechanisms rather than the fact lookup mechanisms needed for robustness.
4 Discussion
Our work evaluates the efficacy of different localization methods for modifying factual associations. We demonstrate clear benefits of localization for editing robustness through localized fine-tuning combined with FLU interpretability.
In Sections 3.3 and 3.4 and Section A.5, we provide evidence that OT and baseline approaches fail to be robust because they target easily-localizable and high direct logit importance extraction components that transform existing latent factual knowledge to the desired output format. This can fail to generalize to different input and output formats and does not target the source of knowledge in the model: other input/output formats can allow alternative mechanisms to extract this knowledge, and relearning can repair the original mechanisms to recover accuracy. In contrast, FLU mechanistic understanding allows us to target editing at the sites where knowledge is sourced, which robustly prevents that information from entering the latent stream in any format.
Acknowledgments
We thank Stephen Casper, Aengus Lynch, Max Li, and Peter Hase for their advice with unlearning methods, interpreting results, and discussing future directions. We also thank Daniel M. Roy, Amr Khalifa, Eleni Triantafillou, Fabien Roger, and Neel Nanda for feedback on various drafts of this work. This project used compute provided by the Center for AI Safety, and part of this project was completed during the ML Alignment & Theory Scholars Program.
Appendix A Appendix
A.1 Sports Unlearning Results
For unlearning on the Sports-Unlearning task, we use a loss function